A spherical balloon is made from a material whose mass is 3.10 kg. The thickness of the material is negligible compared to the 1.30 m radius of the balloon. The balloon is filled with helium (He) at a temperature of 320 K and just floats in air, neither rising nor falling. The density of the surrounding air is 1.19 kg/m³ and the molar mass of helium is 4.0026×10-3 kg/mol. Find the absolute pressure of the helium gas.
Answer #1

Weight of air displaced = weight of balloon + weight of helium

p V g = Mg + mg

Acceleration due to gravity 'g' will cancel on both sides,

1.19 x (4/3) pi x 1.3^3 = 3.1 + m

Mass, m = 7.85 kg

Number of moles, n = 7851.3/4 = 1962.82 moles

Pressure, P = (1962.82 x 8.314 x 320)/(4/3) x pi x 1.3^3 = 5.67 x 10^5 Pa = 5.67 atm

Comment in case any doubt please rate my answer....